# SWIFT: A Monotonic, Flux-Form Semi-Lagrangian Tracer Transport Scheme for Flow with Large Courant Numbers

### v1

@@ 2.1.5 One-Dimensional Scheme for an Advective Tracer
As will be seen in the following sections, the COSMIC splitting uses a one-dimensional advective operator. Although there is flexibility in the choice of this advective operator (for instance, Lin and Rood [1996] use a semi-Lagrangian scheme), in this work we take the approach of Putman and Lin [2007], using conservative operators and the transport of the unity field.
Let $\sigma$ be the unity field, and consider the conservative transport of $\sigma m$. Dividing the conservative transport of $\sigma m$ by the conservative transport of $\sigma$ gives the advective transport of $m$. This means that advective transport of $m$ in the $x$-direction is given by
$m^{n+1}=m^{n}-\Delta{t}\mathcal{A}^{x}(m^{n},u)=\left(m^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},u)\right)/\left(\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u)\right).$ (18)
Following Putman and Lin [2007], this can be rearranged to give an advective increment operator $\mathcal{A}^{x}(m,u)$:
$\Delta{t}\mathcal{A}^{x}(m^{n},u)=m^{n}-\left(m^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},u)\right)/\left(\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u)\right).$ (19)
@@ 2.2 Review of Two-dimensional COSMIC splitting
The notation used will follow that from Sections 2.1.4 and 2.1.5. $\mathcal{F}^{x}(\rho,u^{x})$ represents the one-dimensional flux operator in $x$. The fractional part of this flux is computed using the reconstruction of $\rho$ multiplied by the fractional part of $u^{x}$ (using the methods in Section 2.1), and the integer part being the sum of $\rho$ multiplied by cell volume over the integer number of cells. The one-dimensional flux divergence in $x$ is written as $\nabla_{x}\cdot\mathcal{F}^{x}(\rho,u^{x})$ (similar for $y$). As the fluxes are stored on cell edges, the divergence is at cell centres (co-located with $\rho$). The advective increment operators for $\rho$ are denoted by $\mathcal{A}^{x}(\rho,u^{x})$ and $\mathcal{A}^{y}(\rho,u^{y})$ in $x$ and $y$ respectively (see Section 2.2.2 for the calculation of these increments). Finally, the velocity fields $u^{x}$ and $u^{y}$ are given by (3), with the superscript denoting the orientation of the facet.
@@ 2.2.1 COSMIC Splitting
$\rho^{n+1}=\rho^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\rho_{A}^{y/2},u^{x})-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\rho_{A}^{x/2},u^{y}).$ (21)
The inner half steps are in advective form to preserve constancy, so that a constant $\rho$ subject to a non-divergent flow should remain constant (note that in the literature this constancy is sometimes referred to as consistency – this is distinct from the consistency discussed in our work). This also ensures the stronger requirement that the conservative increments are linear in $\bm{\nabla\cdot u}$. Using $\rho^{x}$ in $\mathcal{F}^{y}$ and $\rho^{y}$ in $\mathcal{F}^{x}$ ensures that all of the cross-terms are captured [Kent, 2019] and that the scheme is symmetric.
@@ 2.2.2 Advective Increment
$\Delta{t}\mathcal{A}^{x}(\rho^{n},u^{x})$ $=\rho^{n}-\left(\rho^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x})\right)/\left(\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x})\right),$ (22a)
$\Delta{t}\mathcal{A}^{y}(\rho^{n},u^{y})$ $=\rho^{n}-\left(\rho^{n}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y})\right)/\left(\sigma-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y})\right),$ (22b)
@@
2.3 Conservative and Consistent Flux-Form Semi-Lagrangian Tracer Transport
This section shows how to achieve consistency between tracers and density for conservative tracer transport using FFSL type schemes and COSMIC splitting.
The notation is expanded as follows. Let $\mathcal{F}^{x}(m,u^{x},\rho)$ represent the one-dimensional flux operator in $x$ (superscript $y$ represents this in $y$). As before, the fractional part of this flux is computed using the reconstruction of $m$ multiplied by the fractional part of $u^{x}$, but now the integer part is the sum of $\rho m$ multiplied by cell volume over the integer number of cells. Note that if the $\rho$ argument is empty, for instance $\mathcal{F}^{x}(m,u^{x})$, then the integer part is simply the sum of $m$ multiplied by cell volume over the integer number of cells. In contrast, when the $\rho$ argument is present, the departure points and fractional flux components need recomputing using that density $\rho$ (see departure point calculation in Section 2.3.1). The flux fields used to transport the density $\rho$ are given by the non-calligraphic $F^{x}$ and $F^{y}$.
2.3.1 One-Dimensional Scheme for a Conservative Tracer
The FFSL schemes are now extended to a tracer mixing ratio $m$ which evolves conservatively relative to a density $\rho$ according to (2). It is assumed throughout that $\rho$ is positive. As discussed in Section 1, the transport of $m$ must be consistent, in the sense that a constant field is preserved. In other words, it is required that the scheme solving for $\rho m$ produces the same results as the scheme solving for $\rho$ when $m=1$. This is achieved by ensuring that if $m=K$ (where $K$ is a constant), the tracer flux $\bm{G}$ is related to the density flux $\bm{F}$ through $\bm{G}=K\bm{F}$. Throughout this section, the tracer flux is written with a $G$ to distinguish it from the density flux $F$.
For Eulerian transport schemes, in order to provide consistency it is common to replace the transporting velocity $\bm{u}$ with the mass flux $\bm{F}$ that was used to transport the density $\rho$. This was the approach used by Bendall et al. [2023] and is essential if $\rho$ and $m$ are not co-located. However, care needs taking to apply this to the FFSL fluxes written in integral form. In the approach presented here, there are two key ingredients to obtain consistency: (a) the computation of departure points specifically for tracer transport (which are in general different from those used for the density transport), and (b) rewriting the integrals as a velocity multiplied by a facet reconstruction (see (15)), which allows the mass flux to be used and consistency to be obtained.
This means that for tracer transport, the integer and fractional components of the Courant number are defined differently to those used for the density transport, although they will coincide when $\rho$ is a constant. The same is true of the fractional part of the flux. To clarify the difference between quantities relating to the tracer quantities and those of the underlying density, the Courant number of tracers is denoted by $k$, while the tracer flux is denoted by $G$. The integer Courant number is determined by:
$k^{int}_{i+1/2}=\begin{cases}\max\left(N\in\mathbb{Z}:|F_{i+1/2}|S_{i+1/2}\Delta t\geq\sum_{j=1}^{N}\rho_{i+1-j}V_{i+1-j}\right),&F_{i+1/2}\geq 0,\\ -\max\left(N\in\mathbb{Z}:|F_{i+1/2}|S_{i+1/2}\Delta t\geq\sum_{j=1}^{N}\rho_{i+j}V_{i+j}\right),&F_{i+1/2}<0,\end{cases}$ (24)
while the fractional component of the mass flux is given by
$\widehat{F}^{frac}_{i+1/2}=\begin{dcases}F_{i+1/2}-\sum_{j=1}^{k^{int}_{i+1/2}}\rho_{i+1-j}V_{i+1-j}/\left(S_{i+1/2}\Delta t\right),&F_{i+1/2}\geq 0,\\ F_{i+1/2}+\sum_{j=1}^{\left|k^{int}_{i+1/2}\right|}\rho_{i+j}V_{i+j}/\left(S_{i+1/2}\Delta t\right),&F_{i+1/2}<0.\end{dcases}$ (25)
It should be emphasised that the flux $\widehat{F}^{frac}$ from (25) is not equal to the fractional part of the mass flux from (15) that was used to transport the density. Finally, the fractional part of the Courant number is given by
$k^{frac}_{i+1/2}=\begin{dcases}\widehat{F}^{frac}_{i+1/2}S_{i+1/2}\Delta t/\left(\rho_{i-k^{int}_{i+1/2}}V_{i-k^{int}_{i+1/2}}\right),&F_{i+1/2}\geq 0,\\ \widehat{F}^{frac}_{i+1/2}S_{i+1/2}\Delta t/\left(\rho_{i+1-k^{int}_{i+1/2}}V_{i+1-k^{int}_{i+1/2}}\right),&F_{i+1/2}<0.\end{dcases}$ (26)
Then the integer components of the tracer flux $G$ is computed as
$G^{int}_{i+1/2}=\begin{dcases}\frac{1}{S_{i+1/2}\Delta{t}}\sum_{j=1}^{k_{i+1/2}^{int}}m_{i+1-j}\rho_{i+1-j}V_{i+1-j},&F_{i+1/2}\geq 0,\\ \frac{1}{S_{i+1/2}\Delta{t}}\sum_{j=1}^{|k_{i+1/2}^{int}|}m_{i+j}\rho_{i+j}V_{i+j},&F_{i+1/2}<0.\end{dcases}$ (27)
If $m=K$, a constant, then with the definition of (24), $G^{int}=K\widehat{F}^{int}=K(F-\widehat{F}^{frac})$.
From Section 2.1 the fractional fluxes for density can be written in the form of (15), the product of the fractional velocity and a facet reconstruction. The same approach can be applied to the fractional component of the tracer flux, so that
$G_{i+1/2}^{frac}=\widehat{F}_{i+1/2}^{frac}\mathcal{R}(m,k_{i+1/2}^{frac}).$ (28)
Now the tracer flux is the product of the density flux and the reconstruction of the tracer mixing ratio. If $m=K$ this yields $G^{frac}=K\widehat{F}^{frac}$, and the tracer transport will be consistent, as $G=G^{frac}+G^{int}=KF$.
$(\rho m)^{n+1}=(\rho m)^{n}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},F^{x},\rho^{n}),$ (29)
where $(\rho m)^{n}\equiv\rho^{n}m^{n}$, and $\mathcal{F}^{x}(m,F^{x},\rho)$ denotes the operator for the evaluation of the fluxes as described in this section, with the fractional part computed using the reconstruction of $m$ multiplied by the fractional part of the density flux $F^{x}$, and the integer part being the sum of $\rho m$ multiplied by the cell volume over the integer number of cells.
@@ 2.3.2 Consistent COSMIC Splitting for Conservative Tracers
The outer step now uses the dry mass fluxes in place of the velocities to compute the tracer fluxes. Following Skamarock [2006], new departure points and fractional mass fluxes need computing, according to (24)-(26). Then the updated mixing ratio is found from
$(\rho m)^{n+1}=(\rho m)^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{y/2},F^{x},\rho^{n})-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(m^{x/2},F^{y},\rho^{n}),$ (31)
@@ 3 Two-dimensional Monotonic SWIFT Splitting
The COSMIC splitting from Sections 2.2 and 2.3.2 provides large time-step-permitting conservative and consistent transport. However, it does not guarantee monotonicity when monotonic limiters are used for the one-dimensional schemes, especially for large Courant number flow (see [Leonard et al., 1996, Lin and Rood, 1996, Bott, 2010] and the results from Section 4).
@@ 3.1 SWIFT Splitting for Density
$\rho_{I}^{x}$ $=\rho^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x}),$ (32a)
$\rho_{I}^{y}$ $=\rho^{n}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y}).$ (32b)
@@
$\sigma^{x}$ $=\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x}),$ (33a)
$\sigma^{y}$ $=\sigma-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y}).$ (33b)
@@
Densities with subscript $I$ denote values obtained from the inner step, while this with subscript $A$ include the division by an appropriately transported unity field to give a density with an advective update. The departure points and fractional components of the wind are recomputed using $\sigma^{y}$ and $\sigma^{x}$ as the density in equations (24)-(26). The outer density step computes the fluxes using reconstructions of $\rho^{x}_{A}$ and $\rho^{y}_{A}$ as in the COSMIC splitting, but now updates the fields from the inner steps:
$\rho^{n+1}=\frac{1}{2}\left[\rho_{I}^{y}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y})+\rho_{I}^{x}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x})\right].$ (35)
@@
$\begin{split}\rho^{n+1}\equiv\rho^{n}&-\frac{\Delta t}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x})-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y})\\ &-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y})-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x}).\end{split}$ (36)
@@
$F^{x}$ $=\frac{1}{2}\left[\mathcal{F}^{x}(\rho^{n},u^{x})+\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y})\right],$ (37a)
$F^{y}$ $=\frac{1}{2}\left[\mathcal{F}^{y}(\rho^{n},u^{y})+\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x})\right],$ (37b)

### v2

@@ 2.1.5 One-Dimensional Scheme for an Advective Tracer
As will be seen in the following sections, the COSMIC splitting uses a one-dimensional advective operator, which corresponds to the discretisation of (1b). Although there is flexibility in the choice of this advective operator (for instance, Lin and Rood [1996] use a semi-Lagrangian scheme), in this work we take the approach of Putman and Lin [2007], using conservative operators and the transport of the unity field.
Let $\sigma$ be the unity field, and consider the conservative transport of $\sigma m$. Dividing the discrete conservative update of $\sigma m$ by the discrete conservative update of $\sigma$ gives a discrete advective update of $m$. This means that the discrete advective update of $m$ in the $x$-direction is given by
$m^{n+1}=m^{n}-\Delta{t}\mathcal{A}^{x}(m^{n},u)=\left(m^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},u,V)\right)/\left(\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u,V)\right).$ (23)
Following Putman and Lin [2007], this can be rearranged to give a discrete operator $\mathcal{A}^{x}(m,u)$ described as the advective increment:
$\Delta{t}\mathcal{A}^{x}(m^{n},u)=m^{n}-\left(m^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},u,V)\right)/\left(\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u,V)\right).$ (24)
2.2 One-Dimensional Flux-Form Semi-Lagrangian Scheme for Conservative Tracer Transport
The one-dimensional FFSL scheme is now extended to a tracer mixing ratio $m$ which evolves conservatively relative to a density $\rho$ according to (2). It is assumed throughout that $\rho$ is positive. As discussed in Section 1, the transport of $m$ must be consistent, in the sense that a constant field is preserved. In other words, it is required that the scheme solving for $\rho m$ produces the same results as the scheme solving for $\rho$ when $m=1$. This is achieved by ensuring that if $m=K$ (where $K$ is a constant), the tracer flux is given by $K\bm{F}$, if the density flux is $\bm{F}$. This property is attained by reusing the mass fluxes that are calculated during the transport of the density variable.
For Eulerian transport schemes, in order to provide consistency it is common to replace the transporting velocity $\bm{u}$ with the mass flux $\bm{F}$ that was used to transport the density $\rho$. This was the approach used by Bendall et al. [2023] and is particularly relevant if $\rho$ and $m$ are not co-located. However, care needs taking to apply this to the FFSL fluxes written in integral form. In the approach presented here, there are two key ingredients to obtain consistency: (a) the computation of departure points specifically for tracer transport (which are in general different from those used for the density transport), and (b) rewriting the integrals as a velocity multiplied by a facet reconstruction (see (15)), which allows the mass flux to be used and consistency to be obtained.
This means that for tracer transport, the integer and fractional components of the Courant number are defined differently to those used for the density transport, although they will coincide when $\rho$ is a constant. The same is true of the fractional part of the flux. However these quantities can all be computed following exactly the same procedure described in Sections 2.1.2 and 2.1.3, but with the following substitutions: $m$ is used in place of $\rho$, $F^{x}$ takes the role of $u^{x}$ and the product $\rho V$ is used instead of just $V$. For the transport of tracers, we can therefore write flux calculations through an operator of the form $\mathcal{F}^{x}(m,F^{x},\rho V)$, whereas the equivalent operator for the transport of $\rho$ is $\mathcal{F}^{x}(\rho,u^{x},V)$. The fractional part of the tracer mass flux is therefore computed using the reconstruction of $m$ multiplied by the fractional part of the density flux $F^{x}$, and the integer part of the flux is the sum of $\rho m$ multiplied by the cell volume over the integer number of cells. With these substitutions, if $m=K$ a constant, the tracer mass flux is $KF^{x}$. This follows directly from the argument at the end of Section 2.1.3.
$(\rho m)^{n+1}=(\rho m)^{n}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}(m^{n},F^{x},\rho^{n}V),$ (26)
where $(\rho m)^{n}\equiv\rho^{n}m^{n}$. The mixing ratio at the next time level is then found from
$m^{n+1}=\frac{(\rho m)^{n}}{\rho^{n}}.$ (27)
As (26) is in flux form, the mass corresponding to $\rho m$ is conserved by this scheme. If $m$ is a constant $K$, then (26) becomes
$(\rho m)^{n+1}=K\rho^{n}-\Delta tK\nabla_{x}\cdot F^{x}.$ (28)
As the corresponding density transport is given by
$\rho^{n+1}=\rho^{n}-\Delta t\nabla_{x}\cdot F^{x},$ (29)
the updated mixing ratio is simply $m^{n+1}=K$ and a constant mixing ratio is preserved. This satisfies the property for consistent tracer transport.
@@ 2.3 Review of Two-dimensional COSMIC splitting
The notation used will follow that from Sections 2.1.4 and 2.1.5 . The $x$ and $y$ variables refer to local directions on a regular quadrilateral mesh, rather than the Cartesian $x$- and $y$-directions. Thus $\mathcal{F}^{x}(\rho,u^{x},V)$ represents the one-dimensional flux operator in the $x$-direction. As the fluxes are stored on cell edges, the divergence is at cell centres (co-located with $\rho$). The advective increment operators for $\rho$ are discretisations of $\bm{u\cdot\nabla}\rho$, and are denoted by $\mathcal{A}^{x}(\rho,u^{x})$ and $\mathcal{A}^{y}(\rho,u^{y})$ in $x$ and $y$ respectively (see Section 2.3.2 for the calculation of these increments). Finally, the velocity fields $u^{x}$ and $u^{y}$ are given by (3) (and so are defined by the mesh, rather than necessarily being physical $x$- and $y$-components), with the superscript denoting the orientation of the facet.
@@ 2.3.1 COSMIC Splitting
$\rho^{n+1}=\rho^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\rho_{A}^{y/2},u^{x},V)-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\rho_{A}^{x/2},u^{y},V).$ (31)
The inner half steps are in advective form which provides the property of divergence-linearity, as if $\rho^{n}=K$, then $\rho_{A}^{x/2}=K$ and $\rho_{A}^{y/2}=K$, which gives $\rho^{n+1}=K(1-\Delta t\bm{\nabla\cdot u})$. Hence a constant $\rho$ subject to a non-divergent flow remains constant. Using $\rho^{x}$ in $\mathcal{F}^{y}$ and $\rho^{y}$ in $\mathcal{F}^{x}$ ensures that all of the cross-terms are captured [Kent, 2019] and that the scheme is symmetric.
@@ 2.3.2 Advective Increment
$\Delta{t}\mathcal{A}^{x}(\rho^{n},u^{x})$ $=\rho^{n}-\left(\rho^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x},V)\right)/\left(\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x},V)\right),$ (32a)
$\Delta{t}\mathcal{A}^{y}(\rho^{n},u^{y})$ $=\rho^{n}-\left(\rho^{n}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y},V)\right)/\left(\sigma-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y},V)\right),$ (32b)
@@ 2.3.3 Consistent COSMIC Splitting for Conservative Tracers
The outer step now uses the dry mass fluxes in place of the velocities to compute the tracer fluxes. Following Skamarock [2006], new departure points and fractional mass fluxes need computing, according to the procedure outlined in Section 2.2. Then the updated mixing ratio is found from
$(\rho m)^{n+1}=(\rho m)^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(m^{y/2},F^{x},\rho^{n}V)-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(m^{x/2},F^{y},\rho^{n}V),$ (34)
@@ 3 Two-dimensional Monotonic SWIFT Splitting
The COSMIC splitting from Section 2.3 provides large time-step-permitting conservative and consistent transport. However, it does not guarantee monotonicity when monotonic limiters are used for the one-dimensional schemes, especially for large Courant number flow (see [Leonard et al., 1996, Lin and Rood, 1996, Bott, 2010] and the results from Section 4).
@@ 3.1 SWIFT Splitting for Density
$\rho_{I}^{x}$ $=\rho^{n}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x},V),$ (35a)
$\rho_{I}^{y}$ $=\rho^{n}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y},V).$ (35b)
@@
$\sigma^{x}$ $=\sigma-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\sigma,u^{x},V),$ (36a)
$\sigma^{y}$ $=\sigma-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\sigma,u^{y},V).$ (36b)
@@
Densities with subscript $I$ denote values obtained from the inner step, while this with subscript $A$ include the division by an appropriately transported unity field to give a density with an advective update. The departure points and fractional components of the wind are recomputed using $\sigma^{y}V$ and $\sigma^{x}V$ as the volume in equations (9)-(11). The outer density step computes the fluxes using reconstructions of $\rho^{x}_{A}$ and $\rho^{y}_{A}$ as in the COSMIC splitting, but now updates the fields from the inner steps:
$\rho^{n+1}=\frac{1}{2}\left[\rho_{I}^{y}-\Delta{t}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y}V)+\rho_{I}^{x}-\Delta{t}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x}V)\right].$ (38)
@@
$\begin{split}\rho^{n+1}\equiv\rho^{n}&-\frac{\Delta t}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x},V)-\frac{\Delta{t}}{2}\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y}V)\\ &-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{n},u^{y},V)-\frac{\Delta{t}}{2}\nabla_{y}\cdot\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x}V).\end{split}$ (39)
@@
$F^{x}$ $=\frac{1}{2}\left[\mathcal{F}^{x}(\rho^{n},u^{x},V)+\mathcal{F}^{x}(\rho^{y}_{A},u^{x},\sigma^{y}V)\right],$ (40a)
$F^{y}$ $=\frac{1}{2}\left[\mathcal{F}^{y}(\rho^{n},u^{y},V)+\mathcal{F}^{y}(\rho^{x}_{A},u^{y},\sigma^{x}V)\right],$ (40b)